2. A simple random sample of 15 people from a certain population has a


mean age of 35 with a standard deviation of 20. Can we conclude that


the mean age of the population is younger than 35? Let alpha = .05.

1
Expert's answer
2022-05-26T17:10:36-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge35"

"H_1:\mu<35"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=14" and the critical value for a left-tailed test is "t_c =-1.76131."

The rejection region for this left-tailed test is "R = \{t:t<-1.76131\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{35-35}{20/\sqrt{15}}=0"

Since it is observed that "t=0>-1.76131=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=14" degrees of freedom, "t=0" is "p=0.5," and since "p=0.5>0.05=\alpha," it is concluded that the null hypothesis is notrejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 35, at the "\alpha = 0.05" significance level.